OpenStudy (anonymous):

A table saw blade, with diameter of 10.00 inches, rotates at about 3200 revolutions per minute. Find the linear speed of a point on the tip of the blade and express it to the nearest mile per hour. A) 71 mph B) 77 mph C) 83 mph D) 89 mph E) 95 mph

OpenStudy (amistre64):

the length it travels in one turn is 10pi inches

OpenStudy (amistre64):

3200(10pi) = 32000 pi inches per minute

OpenStudy (amistre64):

how many inches in a mile?

OpenStudy (amistre64):

abt 5000 ft per mile; little more actually 5000*12 = how many inches in a mile

OpenStudy (amistre64):

72000 inches per mile; 72000 ------- = 72/32pi 32000pi

OpenStudy (amistre64):

a little more than that; but find the actual conversion of inches to miles to get an exact result

OpenStudy (amistre64):

63360 inches per mile -------------------- = ? 32000pi inches

OpenStudy (anonymous):

My conversion looked like this s = 3200 * 2pi * 10 inches t= 1 minute 5280 inches 1 mile --------- = ------- 60 min 1 hour but idk how to do the conversion completely!

OpenStudy (amistre64):

3200 * 2pi5 = 32000pi in / min right?

OpenStudy (amistre64):

63360 inches in one mile ------- 32000pi = 0.63 * 60 minutes i get 37.7 mph

OpenStudy (amistre64):

which aint an option :)

OpenStudy (anonymous):

Ahhhhh Im so confused!

OpenStudy (amistre64):

32000pi is right; 5280*12 = inches per mile

OpenStudy (amistre64):

63360 per mile is right.. 32000pi -------- 63360 maybe?

OpenStudy (amistre64):

that gets me 95 mph :) and thats an option

OpenStudy (anonymous):

My calc was in degrees and not radians, DOY! thank you!

OpenStudy (amistre64):

rad vs deg will tend to mess it up lol